# S2AFormer: Strip Self-Attention for Efficient Vision Transformer

### v1

@@ Abstract
Vision Transformer (ViT) has made significant advancements in computer vision, thanks to its token mixer’s sophisticated ability to capture global dependencies between all tokens. However, the quadratic growth in computational demands as the number of tokens increases limits its practical efficiency. Although recent methods have combined the strengths of convolutions and self-attention to achieve better trade-offs, the expensive pairwise token affinity and complex matrix operations inherent in self-attention remain a bottleneck. To address this challenge, we propose S2AFormer, an efficient Vision Transformer architecture featuring novel Strip Self-Attention (SSA). We design simple yet effective Hybrid Perception Blocks (HPBs) to effectively integrate the local perception capabilities of CNNs with the global context modeling of Transformer’s attention mechanisms. A key innovation of SSA lies in its reducing the spatial dimensions of $K$ and $V$ while compressing the channel dimensions of $Q$ and $K$. This design significantly reduces computational overhead while preserving accuracy, striking an optimal balance between efficiency and effectiveness. We evaluate the robustness and efficiency of S2AFormer through extensive experiments on multiple vision benchmarks, including ImageNet-1k for image classification, ADE20k for semantic segmentation, and COCO for object detection and instance segmentation. Results demonstrate that S2AFormer achieves significant accuracy gains with superior efficiency in both GPU and non-GPU environments, making it a strong candidate for efficient vision Transformers.
Index Terms:
Transformer, strip self-attention, local perception, global context
@@ I Introduction
Vision Transformers (ViTs) [1, 2] have emerged as a notable competitor to CNNs, demonstrating a superior ability to capture long-range interactions between image patches. While Convolutional Neural Networks (CNNs) limit interactions to local regions through shared kernels, ViTs divide the input image into patches and use self-attention (SA) to update token features, enabling global interactions. Consequently, Transformer-based architectures have gained increasing attention in computer vision, achieving remarkable results across various tasks such as image classification [3, 4, 5], object detection [1, 6, 7, 8], and semantic segmentation [9, 10].
[…]
Figure 1: Comparison of our proposed S2AFormer model with SOTA methods (Top-1 accuracy v.s. MACs on ImageNet-1k [18]).
@@
To address computational overhead, various optimized self-attention strategies have been introduced. Methods such as [4, 19, 20] have been developed to limit the self-attention operation to local windows, rather than to the entire feature map. However, this approach compromises global information interaction between patches. Another approach involves token pruning or merging to selectively compute only the most informative tokens and eliminate uninformative tokens. For instance, Dynamic-ViT [21] reduces the token count by eliminating redundant tokens, while EViT [22] merges redundant tokens into a single one. DVT [23] adopts a flexible and dynamic patch-splitting strategy (such as $4\times 4$, $7\times 7$, etc.) tailored to the complexity of each image, moving away from the standard $14\times 14$ approach. MG-ViT [24] emphasizes the need for different regions of an image to receive attention at varying levels of granularity, offering a more nuanced and effective representation. While reducing tokens decreases computation, it presents challenges when dealing with complex, object-rich images, as it becomes increasingly difficult to determine which tokens should be retained.
Some approaches [25, 16, 26] aim to streamline the global sequence length by aggregating tokens across the entire key-value feature space, implementing a coarse-grained global attention mechanism. For instance, the Pyramid Vision Transformer (PVT) [25] uses a large kernel with significant strides to achieve uniform token aggregation, generating a consistent coarse representation across the feature map. Efficient-ViT [27] introduces grouped attention heads to capture multi-scale features while distributing the spatial dimensions of each token, thereby effectively balancing the computational load. Similarly, SG-Former [26] combines the windowing mechanism of the Swin Transformer [4] with the spatial dimensionality reduction strategy of PVT [25], allowing for the extraction of information at different scales. Notably, it further optimizes efficiency by reducing the sequence length of $Q$, significantly decreasing computational complexity. As pointed out in MetaSeg [28], beyond spatial redundancy, significant redundancy also exists in the channel dimension. By leveraging this channel redundancy, the computational complexity can be further reduced. While ViTs excel at modeling long-range dependencies, they tend to struggle with local feature representation compared to CNNs, a limitation that is often overlooked in the aforementioned architectures.
@@
In this paper, we propose an efficient Transformer architecture, named S2AFormer. Our goal is to address the primary challenge of preserving both local sensitivity and global receptive fields while reducing heavy computational overhead. To achieve this, we design simple yet effective Hybrid Perception Blocks (HPBs) in the backbone network. Specifically, we propose a novel Strip Self-Attention (SSA) module that first applies spatial convolutions to the $Key$ and $Value$ features, thereby extracting more distilled feature representations. Subsequently, the $Query$ and $Key$ undergo channel-wise compression, significantly improving memory efficiency and reducing computational complexity from $O(n^{2}\cdot C)$ to $O((\frac{n}{r})^{2}\cdot h)$, where $r$ is the reduction rate and $h$ is the number of attention heads. To compensate for the limited local perceptual capacity of self-attention, we integrate a specialized Local Interaction Module (LIM) into HPB, seamlessly combining global and local receptive fields to improve overall performance. Thus, our method reconciles these complementary strengths through dynamic sparse attention – prioritizing semantically salient regions while filtering redundant spatial correlations – thereby attaining human vision-aligned efficiency without compromising multi-scale representational capacity. This module, composed solely of depth-wise separable and grouped convolutions, further streamlines computational demands. Finally, the integration of SENet [29] further refines the local receptive field by emphasizing salient information. A comprehensive series of rigorous experiments confirms that our S2AFormer architecture either rivals or surpasses the most advanced state-of-the-art (SOTA) backbones (Fig. 1).
@@
    We propose a novel token-mixer mechanism, Strip Self-Attention, performing token interaction in a lightweight manner through both spatial and channel-wise compression, significantly reducing computational cost.
@@
    We introduce Hybrid Perception Blocks (HPBs), a simple yet effective design that enhances the global self-attention with local perceptual capabilities. By incorporating the LIM module, HPBs improve the model’s ability to capture fine-grained visual details, producing more comprehensive and expressive feature representations.
@@
    We pretrain our backbone on ImageNet-1K [18] and evaluate its generalization capability on multiple downstream tasks, including object detection, instance segmentation, and semantic segmentation, all of which demonstrate strong performance, providing a solid pretrained model for future research.
@@
DeiT [5]
PoolFormer [30]
RepViT [31]
SHViT [15]
@@ II-A Efficient ViTs
Originally developed for long-sequence learning in NLP tasks [32], Transformers were later adapted for image classification [2] and for object detection [1], both achieving performance on par with CNNs due to advanced training techniques and larger datasets. DeiT [5] further improved the training process by incorporating distillation, eliminating the need for extensive pretraining [33]. Since then, various modifications of vision Transformers (ViT) and hybrid architectures have emerged, introducing image-specific inductive biases to ViTs, enhancing performance across a range of vision tasks [34, 35, 36, 37, 3, 25]. While ViTs have outperformed CNNs in several vision tasks due to their token mixer’s strong global context capabilities, the complexity of pairwise token interactions and the heavy computational demands of matrix operations present challenges for deploying ViTs in resource-constrained environments and real-time applications.
@@
As a result, researchers have focused on developing various optimization methods to make ViTs more lightweight and better suited for mobile devices. These methods include token pruning or token merging [38, 39, 40, 21], the introduction of novel architectures or modules [41, 42, 43, 44], re-evaluating self-attention and sparse-attention mechanisms [45, 21, 46, 47, 48], and applying search algorithms commonly used in CNNs to identify more compact and faster ViT models [49, 50, 51]. Notable examples include TPS [52], which reduces the number of tokens by truncating based on a threshold. This method shortens the sequence length processed by self-attention and has proved its effectiveness in scenarios where the target objects in the image are relatively simple and singular. EdgeViT [12] improves efficiency by employing a global sparse attention module that focuses on a few selected tokens. In contrast, [25] achieves a better efficiency-accuracy balance by down-sampling the key and value vectors. EdgeNeXt [44] introduces a transposed self-attention mechanism that computes attention maps along the channel dimension rather than the spatial dimension. This approach, combined with token mixing, results in linear complexity relative to the number of tokens. Reformer [53] uses locality-sensitive hashing to group tokens, replacing dot-product attention and reducing complexity from $O(n^{2})$ to $O(nlogn)$. LinFormer [54] employs a low-rank matrix factorization technique to approximate the self-attention matrix, reducing complexity from $O(n^{2})$ to $O(n)$. SwiftFormer [42] introduces a more efficient additive attention mechanism, replacing quadratic matrix multiplications with linear element-wise operations.
@@
While the existing methods can mitigate the quadratic computational complexity of self-attention, they typically achieve this at the expense of accuracy. Furthermore, approaches focusing solely on global self-attention remain inadequate for capturing fine-grained local features, due to the inherent limitations of self-attention in modeling local patterns.
@@ II-B Hybrid Variants
Recent research has introduced a variety of hybrid architectures that integrate CNNs and ViTs within a single framework. By leveraging ViTs’ self-attention for capturing long-range dependencies and CNNs’ local kernels to preserve fine-grained details, these models have achieved enhanced performance across various vision tasks. MobileViT [55] employs a hybrid architecture by combining lightweight MobileNet blocks with multi-head self-attention (MHSA) blocks. The MobileFormer [56] architecture merges MobileNetv3 [57] with ViT [2] to achieve state-of-the-art performance within a 6M-parameter budget, making it well-suited for resource-constrained environments. While these methods achieve fast inference speeds and demonstrate strong suitability for edge deployment, their compromised accuracy presents a critical limitation for performance-sensitive applications.
The CMT [58] architecture introduces a convolutional stem, placing a convolutional layer before each Transformer block, resulting in an alternating sequence of convolutional and Transformer layers. CvT [59] replaces ViTs’ linear embeddings with convolutional token embeddings and uses a convolutional transformer layer to fully leverage these embeddings, ultimately improving performance. However, their improved accuracy comes at the cost of reduced inference speed.
Since 2024, to avoid the computational burden and inference speed bottleneck caused by self-attention, efficient methods have been exploring the use of convolution-based techniques as alternatives in the token mixer component. For example, RepViT [31] builds upon MobileNet-V3 [57] by streamlining the block structure and removing the residual connection used during training to achieve faster inference speeds. RepNeXt [60] utilizes multi-scale feature representations while combining both serial and parallel structural reparameterization (SRP) methods to expand the network’s depth and width, resulting in a unique convolutional attention mechanism. CAS-ViT [61] adopts a serial structure with CBAM [62], where spatial and channel-wise weights for the $q$ and $k$ are calculated. These weights are then summed and used to perform matrix calculations with $v$, resulting in an additive attention mechanism. This is a promising approach, given that the computational complexity of convolution is $O(n)$. However, purely CNN-based architectures that rely on large-kernel convolutions still fall short in performance compared to self-attention mechanisms, despite offering notable improvements in speed and efficiency.
@@
Figure 3: Overview of our proposed S2AFormer. Similar to [4, 25], we employ a hierarchical architecture with four stages, each containing $L_{i}$ Hybrid Perception Blocks. Table II provides the detailed network configurations of S2AFormer variants.
@@
Model Blocks $L$ Channels $C$Para. (M) GMACs
@@ III-A Overall Architecture
Building on these insights, we propose S2AFormer, an efficient Transformer-based backbone featuring a hybrid architecture that effectively preserves both local and global receptive fields while significantly reducing computational overhead. The overall structure of S2AFormer, as illustrated in Fig. 3, adopts the four-stage framework commonly used in traditional CNNs [63] and hierarchical ViTs [4], enabling the generation of multi-scale feature maps suitable for dense prediction tasks such as object detection and semantic segmentation.
Our approach begins with a Stem Block consisting of three $3\times 3$ convolutional layers, with the first layer using a stride of 2 [64, 37]. This is followed by four stages, each responsible for generating feature maps at different scales. These stages share a consistent architecture, featuring a patch embedding and multiple $L_{i}$ Hybrid Perception Blocks (HPBs). The HPBs, equipped with two core modules, SSA and LIM, effectively capture both local and global dependencies without the heavy computation overhead.
@@
III-B Hybrid Perception Block
The HPB framework, shown in Fig. 3(c), consists of two main components: the Strip Self-Attention (SSA) module in Fig. 3(d) and the Local Interaction Module (LIM) in Fig. 3(e). The overall operation is defined as follows:
$f_{Conv}=DWConv\left(x\right)+x,$ (1)
$f_{ssa}=SSA\left({LN\left(f_{Conv}\right)}\right)+f_{Conv},$ (2)
$f_{lim}=LIM\left({LN\left(f_{ssa}\right)}\right)+f_{ssa}.$ (3)
Finally, the MLP is defined as:
$f_{mlp}=MLP\left({LN\left(f_{lim}\right)}\right)+f_{lim}.$ (4)
III-B1 Strip Self-Attention (SSA)
In the standard self-attention mechanism [2], the input $x\in{\mathbb{R}^{N\times C}}$, where $N=HW$, is linearly projected into the query $Q\in{\mathbb{R}^{N\times d_{q}}}$, the key $K\in{\mathbb{R}^{N\times d_{k}}}$, and the value $V\in{\mathbb{R}^{N\times d_{v}}}$. Subsequently, the self-attention mechanism is executed as follows:
To reduce computational complexity, a Convolution Spatial Reduction (CSR) [25] is employed, which uses a $k\times k$ depth-wise convolution with a stride of $k$ to downsample the spatial dimensions of $x$ before the linear projection operation. This results in $\widetilde{x}\in{\mathbb{R}^{N_{s}\times C}}$, where $N_{s}=\frac{N}{{{k^{2}}}}$. In our work, we propose a Strip Self-Attention (SSA) module as an innovative token mixing mechanism in the HPB to balance global feature extraction and the computational efficiency of self-attention.
As shown in Fig. 3(d), the channel dimensions of the query $Q$ and key $K$ are compressed into a single dimension to further minimize computational overhead. Experimental results demonstrate that a strip-shaped configuration for $Q\in{\mathbb{R}^{N\times 1}}$ and $K\in{\mathbb{R}^{N_{s}\times 1}}$ can effectively capture global similarities. The pseudocode for the entire operation of SSA is presented in Algorithm 1, followed by a detailed complexity analysis.
III-B2 Local Interaction Module (LIM)
where $SE$ denotes the channel attention module [29].
@@ III-B2 Local Interaction Module (LIM)
         $\widetilde{x}\leftarrow\text{Spatial Reduction}(x)$
         $K\leftarrow\text{Channel Squeeze}(\text{Linear Projection}(\widetilde{x}))$
         $V\leftarrow\text{Linear Projection}(\widetilde{x})$

### v2

@@ Abstract
The Vision Transformer (ViT) has achieved remarkable success in computer vision due to its powerful token mixer, which effectively captures global dependencies among all tokens. However, the quadratic complexity of standard self-attention with respect to the number of tokens severely hampers its computational efficiency in practical deployment. Although recent hybrid approaches have sought to combine the strengths of convolutions and self-attention to improve the performance–efficiency trade-off, the costly pairwise token interactions and heavy matrix operations in conventional self-attention remain a critical bottleneck. To overcome this limitation, we introduce S2AFormer, an efficient Vision Transformer architecture built around a novel Strip Self-Attention (SSA) mechanism. Our design incorporates lightweight yet effective Hybrid Perception Blocks (HPBs) that seamlessly fuse the local inductive biases of CNNs with the global modeling capability of Transformer-style attention. The core innovation of SSA lies in simultaneously reducing the spatial resolution of the key ($K$) and value ($V$) tensors while compressing the channel dimension of the query ($Q$) and key ($K$) tensors. This joint spatial-and-channel compression dramatically lowers computational cost without sacrificing representational power, achieving an excellent balance between accuracy and efficiency. We extensively evaluate S2AFormer on a wide range of vision tasks, including image classification (ImageNet-1K), semantic segmentation (ADE20K), and object detection/instance segmentation (COCO). Experimental results consistently show that S2AFormer delivers substantial accuracy improvements together with superior inference speed and throughput across both GPU and non-GPU platforms, establishing it as a highly competitive solution in the landscape of efficient Vision Transformers.
@@ I Introduction
Vision Transformers (ViTs) [carion2020end, dosovitskiy2020image] have emerged as a notable competitor to CNNs, demonstrating a superior ability to capture long-range interactions between image patches. While Convolutional Neural Networks (CNNs) limit interactions to local regions through shared kernels, ViTs divide the input image into patches and use self-attention (SA) to update token features, enabling global interactions. Consequently, Transformer-based architectures have gained increasing attention in computer vision, achieving remarkable results across various tasks such as image classification [liu2021swin, touvron2021training], object detection [carion2020end, maaz2022class], and semantic segmentation [cheng2022masked, xie2021segformer].
Figure 1: Comparison of our proposed S2AFormer model with SOTA methods, including EMOv2 [zhang2025emov2], SwiftFormer [shaker2023swiftformer], and iFormer [zheng2025iformer]. (Top-1 accuracy v.s. MACs on ImageNet-1k [deng2009imagenet]). In theory, the optimal performance is in the upper-left of the plot, which means higher top-1 accuracy and fewer MACs.
@@
To address computational overhead, various optimized self-attention strategies have been introduced. Methods such as [liu2021swin, dong2022cswin] have been developed to limit the self-attention operation to local windows, rather than to the entire feature map. However, this approach compromises global information interaction between patches. Another approach involves token pruning or merging to selectively compute only the most informative tokens and eliminate uninformative tokens. For instance, Dynamic-ViT [rao2021dynamicvit] reduces the number of tokens by eliminating redundant tokens, while EViT [liang2022not] merges redundant tokens into a single token. DVT [wang2021not] adopts a flexible and dynamic patch-splitting strategy (e.g., $4\times 4$, $7\times 7$, etc.) tailored to the complexity of each image, moving away from the standard $14\times 14$ approach. MG-ViT [zhang2024mg] emphasizes the need for different regions of an image to receive attention at varying levels of granularity, offering a more nuanced and effective representation. While reducing tokens decreases computation, it presents challenges when dealing with complex, object-rich images, as it becomes increasingly difficult to determine which tokens should be retained.
Some approaches [wang2021pyramid, wang2021pvtv2, ren2023sg] aim to streamline the global sequence length by aggregating tokens across the entire key-value feature space, implementing a coarse-grained global attention mechanism. For instance, the Pyramid Vision Transformer (PVT) [wang2021pyramid] uses a large kernel with significant strides to achieve uniform token aggregation, generating a consistent coarse representation across the feature map. Efficient-ViT [xie2023efficient] introduces grouped attention heads to capture multi-scale features while distributing the spatial dimensions of each token, thereby effectively balancing the computational load. Similarly, SG-Former [ren2023sg] combines the windowing mechanism of the Swin Transformer [liu2021swin] with the spatial dimensionality reduction strategy of PVT [wang2021pyramid], allowing for the extraction of information at different scales. Notably, it further optimizes efficiency by reducing the sequence length of $Q$, significantly decreasing computational complexity. As pointed out in MetaSeg [kang2024metaseg], beyond spatial redundancy, significant redundancy also exists in the channel dimension. By leveraging this channel redundancy, the computational complexity can be further reduced. While ViTs excel at modeling long-range dependencies, they tend to struggle with local feature representation compared to CNNs, a limitation that is often overlooked in the aforementioned architectures.
@@
In this paper, we propose an efficient Transformer architecture, named S2AFormer. Our goal is to address the primary challenge of preserving both local sensitivity and global receptive fields while reducing heavy computational load. To achieve this, we first design simple yet effective Hybrid Perception Blocks (HPBs) in the backbone network. Specifically, we propose a novel Strip Self-Attention (SSA) module that first applies spatial convolutions to the $Key$ and $Value$ features, thereby extracting more distilled feature representations. Subsequently, the $Query$ and $Key$ undergo channel-wise compression, significantly improving memory efficiency and reducing computational complexity from $O(n^{2}\cdot C)$ to $O((\frac{n}{r})^{2}\cdot h)$, where $r$ is the reduction rate and $h$ is the number of attention heads.
@@
Meanwhile, rotations and translations are common data augmentations in vision tasks, and a well-designed model should ideally produce consistent predictions under such transformations, demonstrating a desirable degree of geometric invariance. However, Transformers relying on absolute positional encodings often introduce patch-specific biases that anchor predictions to fixed spatial coordinates, thereby undermining this invariance. Moreover, purely global attention mixing tends to neglect fine-grained local neighborhood structures and precise boundary details within patches. To address these limitations, we incorporate a dedicated Local Interaction Module (LIM) into HPB. This module effectively complements global receptive fields with localized processing, significantly enhancing boundary fidelity and robustness to rotations and translations. The LIM is implemented using only lightweight depthwise separable and grouped convolutions, ensuring minimal additional computational overhead. Furthermore, we integrate SENet [hu2018squeeze] to dynamically reweight features, enabling the model to focus on semantically salient information within the local receptive field. A comprehensive series of rigorous experiments confirms that our S2AFormer architecture either rivals or surpasses the most advanced state-of-the-art backbones (as shown in Fig. 1).
@@
    We propose a novel token-mixer mechanism, termed SSA, which enables token interaction in a lightweight manner through both spatial and channel-wise compression, effectively reducing computational cost and achieving faster inference.
@@
    We introduce HPB, a simple yet effective design that enhances the global self-attention with local perceptual capabilities. By incorporating the LIM module, HPB improves the model’s ability to capture fine-grained visual details, producing more comprehensive and expressive feature representations.
@@
    We pretrain our backbone on ImageNet-1K [deng2009imagenet] and evaluate its generalization capability on multiple downstream tasks, including object detection, instance segmentation, and semantic segmentation, all of which demonstrate strong performance, providing a solid pretrained model for future research.
@@
DeiT [touvron2021training]
PoolFormer [yu2022metaformer]
RepViT [wang2024repvit]
SHViT [yun2024shvit]
@@ II-A Efficient ViTs
Originally developed for long-sequence learning in NLP tasks [vaswani2017attention], Transformers were later adapted for image classification [dosovitskiy2020image] and for object detection [carion2020end], both achieving performance on par with CNNs due to advanced training techniques and larger datasets. DeiT [touvron2021training] further improved the training process by incorporating distillation, eliminating the need for extensive pretraining [yuan2021tokens]. Since then, various modifications of vision Transformers (ViT) and hybrid architectures have emerged, introducing image-specific inductive biases to ViTs, enhancing performance across a range of vision tasks [xu2024haformer, wang2021pyramid]. While ViTs have outperformed CNNs in several vision tasks due to their token mixer’s strong global context capabilities, the complexity of pairwise token interactions and the heavy computational demands of matrix operations present challenges for deploying ViTs in resource-constrained environments and real-time applications.
@@
As a result, researchers have focused on developing various optimization methods to make ViTs more lightweight and better suited for mobile devices. These methods include token pruning or token merging [kim2024token, tang2023dynamic, rao2021dynamicvit], the introduction of novel architectures or modules [shaker2023swiftformer, li2022efficientformer, maaz2022edgenext], re-evaluating self-attention and sparse-attention mechanisms [fan2025breaking, ali2021xcit, ma2024efficient], and applying search algorithms commonly used in CNNs to identify more compact and faster ViT models [zhou2022training]. Notable examples include TPS [wei2023joint], which reduces the number of tokens by truncating based on a threshold. This method shortens the sequence length processed by self-attention and has proved its effectiveness in scenarios where the target objects in the image are relatively simple and singular. EdgeViT [chen2022edgevit] improved efficiency by employing a global sparse attention module that focuses on a few selected tokens. In contrast, [wang2021pyramid] achieves a better efficiency-accuracy balance by down-sampling the key and value vectors. EdgeNeXt [maaz2022edgenext] introduced a transposed self-attention mechanism that computes attention maps along the channel dimension rather than the spatial dimension. This approach, combined with token mixing, results in linear complexity relative to the number of tokens. Reformer [kitaev2020reformer] used locality-sensitive hashing to group tokens, replacing dot-product attention and reducing complexity from $O(n^{2})$ to $O(nlogn)$. LinFormer [wang2020linformer] employed a low-rank matrix factorization technique to approximate the self-attention matrix, reducing complexity from $O(n^{2})$ to $O(n)$. Likewise, RAVLT [fan2025breaking] introduced a rank-augmented linear attention framework that effectively balances the performance of traditional softmax attention with the efficiency of linear attention methods. SwiftFormer [shaker2023swiftformer] introduces a more efficient additive attention mechanism, replacing quadratic matrix multiplications with linear element-wise operations.
@@
Although existing methods can alleviate the quadratic computational complexity of self-attention, they often do so at the cost of reduced accuracy. Moreover, approaches that rely exclusively on global self-attention remain insufficient for capturing fine-grained local features, owing to the intrinsic limitations of self-attention in modeling detailed local patterns.
Figure 3: Overview of our proposed S2AFormer. Similar to [liu2021swin, wang2021pyramid], we employ a hierarchical architecture with four stages, each containing $L_{i}$ Hybrid Perception Blocks. Table II provides the detailed network configurations of S2AFormer variants.
@@ II-B Hybrid Variants
Recent research has introduced a variety of hybrid architectures that integrate CNNs and ViTs within a single framework. By leveraging ViTs’ self-attention for capturing long-range dependencies and CNNs’ local kernels to preserve fine-grained details, these models have achieved enhanced performance across various vision tasks. MobileViT [mehta2021mobilevit] employed a hybrid architecture by combining lightweight MobileNet blocks with multi-head self-attention (MHSA) blocks. The MobileFormer [chen2022mobile] architecture merged MobileNetv3 [koonce2021mobilenetv3] with ViT [dosovitskiy2020image] to achieve state-of-the-art performance within a 6M-parameter budget, making it well-suited for resource-constrained environments. While these methods achieve fast inference speeds and demonstrate strong suitability for edge deployment, their compromised accuracy presents a critical limitation for performance-sensitive applications.
The CMT [guo2022cmt] architecture introduced a convolutional stem, placing a convolutional layer before each Transformer block, resulting in an alternating sequence of convolutional and Transformer layers. CvT [wu2021cvt] replaced ViTs’ linear embeddings with convolutional token embeddings and uses a convolutional transformer layer to fully leverage these embeddings, ultimately improving performance. However, their improved accuracy comes at the cost of reduced inference speed.
Since 2024, to avoid the computational burden and inference speed bottleneck caused by self-attention, efficient methods have been exploring the use of convolution-based techniques as alternatives in the token mixer component. For example, RepViT [wang2024repvit] built upon MobileNet-V3 [koonce2021mobilenetv3] by streamlining the block structure and removing the residual connection used during training to achieve faster inference speeds. RepNeXt [zhao2024repnext] utilized multi-scale feature representations while combining both serial and parallel structural reparameterization (SRP) methods to expand the network’s depth and width, resulting in a unique convolutional attention mechanism. CAS-ViT [zhang2024cas] adopted a serial structure with CBAM [woo2018cbam], where spatial and channel-wise weights for the query ($Q$) and key ($K$) are calculated. These weights are then summed and used to perform matrix calculations with value ($V$), resulting in an additive attention mechanism. This is a promising approach, given that the computational complexity of convolution is $O(n)$. However, purely CNN-based architectures that rely on large-kernel convolutions still fall short in performance compared to self-attention mechanisms, despite offering notable improvements in speed and efficiency.
@@
Model Blocks $L$ Channels $C$ #Para. (M) GMACs
@@
L [4,4,20,4] [96, 192, 384, 768] 76.58 12.53
Figure 4: Comparison of different self-attention mechanisms: (a) Vanilla self-attention in ViTs [dosovitskiy2020image], which computes global attention using standard dot-product operations. (b) Separable self-attention in MobileViT-v2 [mehta2022separable], which applies element-wise operations on query ($Q$) and key ($K$) to form a context vector. (c) Swift self-attention in SwiftFormer [shaker2023swiftformer], where $Q$ is weighted and pooled into global queries, broadcast, and multiplied element-wise with $K$ to generate global context. (d) Convolutional additive self-attention in CAS-ViT [zhang2024cas], which replaces the global dot-product with a cascaded design applying spatial attention followed by channel attention. (e) Our proposed strip self-attention jointly compresses spatial and channel dimensions to effectively eliminate redundant information, achieving a lightweight design while preserving dense global dependencies.
@@ III-A Overall Architecture
Building on these insights, we propose S2AFormer, an efficient Transformer-based backbone featuring a hybrid architecture that effectively preserves both local and global receptive fields while significantly reducing computational overhead. The overall structure of S2AFormer, as illustrated in Fig. 3, adopts the four-stage framework commonly used in traditional CNNs [he2016deep] and hierarchical ViTs [liu2021swin], enabling the generation of multi-scale feature maps suitable for dense prediction tasks such as object detection and semantic segmentation.
Our approach begins with a Stem Block consisting of three $3\times 3$ convolutional layers, with the first layer using a stride of 2 [li2023rethinking]. This is followed by four stages, each responsible for generating feature maps at different scales. These stages share a consistent architecture, featuring a patch embedding and multiple $L_{i}$ Hybrid Perception Blocks (HPBs). The HPBs, equipped with two core modules, SSA and LIM, effectively capture both local and global dependencies without the heavy computation overhead.
@@
  if $\text{sr\_ratio}>1$ then
   $\widetilde{x}\leftarrow\text{Spatial Reduction}(x)$
   $)K\leftarrow\text{Channel Squeeze}(\text{Linear Projection}(\widetilde{x}$))